Please help me find the value of each in the space provided, showing all steps.
Given: f(x)=3x-7 G(x)=x^2+3
task 5: Find f(g(2))


thanks

Answer by checkley77(12844)

Find f(g(2))
f(x)=3x-7
g(x)=x^2+3
g(2)=2^2+3=4+3=7
f(g(2)=f(7)
f(7)=3*7-7=21-7=14 ans.